A firm evaluates all of its projects by using the NPV decision rule.

Year Cash Flow

0 –$27,000

1 22,000

2 14,000

3 7,000

a. At a required return of 28 percent, what is the NPV for this project?

b. At a required return of 39 percent, what is the NPV for this project?
Answer #1

a.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=22,000/1.28+14,000/1.28^2+7,000/1.28^3

=29070.28

NPV=Present value of inflows-Present value of outflows

=29070.28-27,000

=$2070.28(Approx).

b.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=22,000/1.39+14,000/1.39^2+7,000/1.39^3

=25679.82

NPV=Present value of inflows-Present value of outflows

=25679.82-27,000

=$(1320.18)(Approx).(Negative).